Average: 1S 5. for negative externalities- Regulation versus tradable permits Suppose the government wants to reduce the total pollution emitted by three local firms. Currently, each firm is creating 4 units of pollution in the area, for a total of 12 pollution units. If the government wants to reduce total pollution in the area to 6 units, it can choose between the following two methods: Available Methods to Reduce Pollution 1. The government sets pollution standards using regulation 2. Th e government all Each firm faces different costs, so reducing pollution is more difficult for some firms than others. The following table shows the cost each firm faces to eliminate each unit of pollution. For each firm, assume that the cost of reducing pollution to zero (that is, eliminating alil 4 units of pollution is prohibitively expensive. Cost of Eliminating the... Second Unit of Pollution Dollars) 125 70 800 First Unit of Pollution (Dollars) 90 Third Unit of Pollution (Dollars) Firm Firm X Firm Y Firm Z 180 110 1,500 650
media%2Fd68%2Fd688b954-74ad-499a-9982-1f
media%2F20b%2F20bb3a59-6ebe-4303-a845-fc
media%2F2a7%2F2a79652b-5026-4838-9045-56
media%2F175%2F175b82f6-2324-40b6-bed7-fa
media%2Fe1f%2Fe1f78652-906d-4156-9b71-4b
Answer #1

Ans:

1) The total cost of eliminating 2 units is pollution

firm X = 90+ 125

= $215

firm Y = 55+70

= $125

firm Z = 650+800

= $1450

2) $579,$787

The cost of eliminating 3rd unit of pollution for firm X and firm Y is $180 and $110.Hence at a price which is above $180 both firm X and firm Y will sell a permit to firm Z.

3) when the permit price is $498 per permit,

Firm Initial pollution permit allocation Action Final amount of pollution eliminated cost of pollution reduction
X 2 sell one permit 3 units $90 + $125 + $180 =$395
Y 2 sell one permit 3 units $55 + $70 + $110 =$235
Z 2 buy two permits 0 units 0

4)

The total cost of eliminating six units of pollution

Regulation = $215 + $125 + $1450

= $1790

Tradable permits = $395 + $235

= $630

5) Less

eliminating pollution using tradable permits is less costly to society than regulation.

> for number 5 my correct answer was more not less

MATTHEW URIBE Tue, Nov 9, 2021 1:49 PM